### v1

# Enhancing Transport Barriers with Swimming Microorganisms in Chaotic Flows

Abstract
We investigate the effects of bacterial activity on the mixing and transport properties of two-dimensional, time-periodic flows in experiments and in a simple model. We focus on the interactions between swimming E. coli and the flow Lagrangian Coherent Structure (LCS), which are computed from experimentally measured velocity fields. Experiments show that such interactions are non-trivial and can lead to transport barriers through which the tracer flux is significantly reduced. Using the Poincaré map, we show that these transport barriers coincide with the outermost members of elliptic LCSs known as Lagrangian vortex boundaries. Numerical simulations further show that elliptic LCSs can repel elongated swimmers and lead to swimmer depletion within Lagrangian coherent vortices. A simple mechanism shows that such depletion is due to the preferential alignment of elongated swimmers with the tangents of elliptic LCSs. Our results provide insights into understanding the transport of microorganisms in complex flows with dynamical topological features from a Lagrangian viewpoint.
keywords:
1 Introduction
Many microorganisms live in environments characterized by currents (e.g., oceans, rivers, human intestines), and flow mediates many important microbial processes such as infection (Costerton et al., 1999; Yawata et al., 2016), nutrients uptake (Taylor & Stocker, 2012; Musielak et al., 2009), and reproduction (Riffell & Zimmer, 2007; Zimmer & Riffell, 2011). Flow exerts forces and torques on microorganisms that affect their motility dynamics and spatial distribution (Guasto et al., 2012; Wheeler et al., 2019). It also controls the transport of essential molecules including nutrients, oxygen, and signals of mates and predators (Kim et al., 2016). Microorganisms, in turn, can adapt their swimming behavior to these physical and chemical gradients (Stocker et al., 2008). Such interactions can lead to non-trivial phenomena such as rheotaxis (Hill et al., 2007; Marcos et al., 2012; Mathijssen et al., 2019), gyrotaxis (De Lillo et al., 2014; Borgnino et al., 2018), and chemotaxis (Locsei & Pedley, 2009; Stocker & Seymour, 2012), to name a few examples.
Even simple shear flows, when coupled to cell motility and morphology, can give rise to complex transport and cell motility behavior (Rusconi et al., 2014; Durham et al., 2009; Ezhilan & Saintillan, 2015; Gustavsson et al., 2016). For instance, fluid shear can cause a torque that constantly rotates elongated cell body and can result in accumulation of motile bacteria and phytoplankton in high-shear-rate regions of the flow (Rusconi et al., 2014; Barry et al., 2015; Ezhilan & Saintillan, 2015). Near surfaces, shear flows can orient flagellated cells against the flow, causing bacteria and spermatozoa to swimming upstream (Tung et al., 2015; Zaferani et al., 2018; Mathijssen et al., 2019). Away from surfaces, shear gradients can trap bottom-heavy gyrotactic swimmers at certain depth of water column, causing the formation of intense cell assemblages called “thin layers” (Durham et al., 2009, 2013; Gustavsson et al., 2016). In unsteady and/or complex flows, transport of microorganisms shows intriguing phenomena such as aggregation, dispersion, and pattern formation (Torney & Neufeld, 2007; Khurana et al., 2011; Khurana & Ouellette, 2012; Zhan et al., 2014; Qin & Arratia, 2022), but are less understood. In time-periodic flows, simulations show that elongated swimmers can be trapped or repelled by elliptic islands depending on their shape and swimming speed (Torney & Neufeld, 2007); such trapping effects can lead to a reduction in long-term swimmer transport (Khurana et al., 2011). Recent experiments and simulations show that microswimmers can be trapped, repelled, or dispersed by vortices depending on the dimensionless path length and swimming speed (Qin & Arratia, 2022). In isotropic turbulence, simulations show that elongated swimmers, while remaining rather uniformly distributed, exhibit preferential alignment with instantaneous Eulerian fields such as local velocity (Borgnino et al., 2019), vorticity (Zhan et al., 2014) and velocity gradient (Pujara et al., 2018).
Recently, it has been shown that Lagrangian Coherent Structure (LCS) can be a useful concept to understand the transport properties of swimming microorganisms in complex flows in both numerical simulations and experiments (Khurana & Ouellette, 2012; Dehkharghani et al., 2019; Ran et al., 2021; Si & Fang, 2021, 2022; Yoest et al., 2022). Simulations in chaotic flows show that elongated swimmers align with repelling LCSs of hyperbolic fixed points (Khurana & Ouellette, 2012), while later numerical studies show that elongated active particles have a much stronger alignment with attracting LCSs (Si & Fang, 2021, 2022), similar to passive elongated particles (Parsa et al., 2011). Experimental investigations that examine the interactions of swimming organisms and flow LCSs are few but show some intriguing phenomena. Experiments in model porous media show that bacteria align and accumulate near attracting LCSs and induce filamentous density patters (Dehkharghani et al., 2019). In time-periodic flows, experiments show that the accumulation of bacteria near attracting LCSs can attenuate stretching and hinder large-scale transport, although small-scale mixing is locally enhanced (Ran et al., 2021). Most, if not all, previous studies focus on attracting and repelling LCSs associated with the flow hyperbolic fixed points. That is not surprising since one expects large levels of strain near or around those fixed points. Less understood are swimmer interactions with elliptic LCSs, i.e., vortex-like flow dynamical features (Haller, 2015; Haller et al., 2016; Farazmand & Haller, 2016). That is the focus of this manuscript.
Here, we experimentally investigate the effects of bacterial activity on the mixing and transport properties of a passive scalar in a time-periodic flow in experiments and simulations. We focus on the interaction of swimming bacteria (E. coli) with the elliptic LCSs of the flow. Results show that such interaction leads to transport barriers through which the fluxes of the passive tracer are significantly reduced. By constructing the Poincaré map from velocimetry data, we show that these transport barriers coincide with the outermost member of elliptic LCSs, or namely, Lagrangian vortex boundaries (LVBs). We further test these results in numerical simulations and find that elliptic LCSs repel elongated swimmers and lead to swimmer accumulation outside (or swimmer depletion inside) the Lagrangian vortices. A simple mechanism shows that the repulsion of swimmers is due to the preferential alignment of elongated swimmers with the tangents of elliptic LCSs. Overall, these results can be useful in understanding the transport of microorganisms in chaotic flows with elliptic dynamical features or nontrivial vortex structures.
[…]
2 Methods
[…]
3 Results and Discussion
3.1 Experimental Results
Figures 1(a) and 1(b) show sample snapshots of dye mixing in the buffer solution ($\phi_{b}=0\%$) and active suspension ($\phi_{b}=0.5\%$), respectively. Both snapshots are taken after $N=300$ periods of forcing, with dye initially confined on the right half of the images at $N=0$. The dye patterns reveal complex patterns and the underlying vortex structures, which are similar in the buffer and the active suspension. However, a main difference is the existence of undyed (void-like) regions near the center of vortex structures in the active suspension [Fig. 1(b)], while similar regions are dyed in the buffer case [Fig. 1(a)]. The undyed void-like regions near the vortex center appear primarily in the left half of the image that is initially not covered by dye. This suggests that the interplay of microbial activity and vortex structures leads to (enhanced) transport barriers through which the dye fails to penetrate or penetrates much more slowly. A video of the dye mixing processes in the buffer and the active suspension are shown in Movie 1.
To further understand the interaction between activity and vortex structures, we plot in Fig. 1(d) the flow Poincaré map obtained from PTV experiments. Lines in the map are stroboscopic particle trajectories that connect a particle’s initial position to its next position after a period ($T=5$ s). These trajectories are obtained by numerical integration of fluid particles in the experimentally measured velocity fields. The Poincaré map reveals several nested families of closed trajectories, whose behaviors mimic Kolmogorov–Arnold–Moser (KAM) tori in Hamiltonian dynamical systems (Ottino, 1989). These nested torus families define Lagrangian coherent vortices known as the elliptic LCSs, with each torus acting as a barrier that blocks tracer (dye) transport within the vortex (Katsanoulis et al., 2020; Haller et al., 2020; Aksamit & Haller, 2022). Unlike hyperbolic LCSs, elliptic LCSs are rotation-dominated regions that do not experience substantial (fluid) stretching. We note that Lagrangian coherent vortices are fundamentally different from the vortices defined by Eulerian fields such as vorticity. While we find a striking correspondence between the dye mixing patterns [Figs. 1(a) and 1(b)] and the Lagrangian coherent vortices in the Poincaré map [Fig. 1(d)], such similarity is absent for the Eulerian vorticity field in Fig. 1(c).
To quantitatively locate the elliptic LCSs, we calculate the trajectory rotation average (TRA) in the Poincaré map, as shown by the color code in Fig. 1(d). The TRA characterizes the average angular velocity of a particle trajectory and is defined as:
$\overline{\mathrm{TRA}}_{0}^{N}(\mathbf{x}_{0})=\frac{1}{NT}\sum_{i=0}^{N-1}\cos^{-1}\frac{\langle\dot{\mathbf{x}}_{i},\dot{\mathbf{x}}_{i+1}\rangle}{\left|\dot{\mathbf{x}}_{i}\right|\left|\dot{\mathbf{x}}_{i+1}\right|}.$ (1)
Here, $\dot{\mathbf{x}}_{i}$ is a stroboscopic velocity related to the rate of change of a particle’s position in the Poincaré map, and $N$ is the number of periods over which the quantity is calculated. Although the TRA is originally defined for actual particle trajectories (Haller et al., 2021), we use it here for the stroboscopic trajectories in the extended phase space of the Poincaré map. We interpolate the values of the TRA onto a uniform spatial grid to obtain the TRA field (see Movie 2). In two-dimensional flows, elliptic LCSs can be located from the closed convex contours of the TRA field (Haller et al., 2021; Aksamit & Haller, 2022). Since elliptic LCSs are nested families, we identify the outermost convex TRA contours as the boundaries of Lagrangian coherent vortices, and the centroid of the innermost convex TRA contours as elliptic fixed points. For concision, we refer to the outermost members of elliptic LCSs as Lagrangian vortex boundaries (LVBs).
Figure 2: Enlarged photographs of the dye concentration field for (a) buffer and (b) active suspension, in the dashed square regions shown in Fig. 1. Green contours are Lagrangian vortex boundaries (LVBs) identified from the TRA field and red dots are elliptic fixed points. (c) Dye flux as a function of time $N$, for the vortex with the label “4”. Inset: average dye flux as a function of vortex Reynolds number $Re_{\Gamma}$, for the buffer (blue) and the active suspension (red). Different markers represent different vortices in the flow field. Error bars are standard deviations.
Figures 2(a) and 2(b) show Lagrangian vortex boundaries (LVBs) (green contours) and elliptic fixed point (red dots) identified from the TRA field, superimposed on the dye mixing patterns in the buffer and the active suspension, respectively. Although LVBs in chaotic flows are known to correspond to transport barriers (Katsanoulis et al., 2020; Haller et al., 2020; Aksamit & Haller, 2022), the LVBs in the buffer case [Fig. 2(a)] contain dye inside. On the other hand, the LVBs in the active case [Fig. 2(b)] enclose void-like undyed regions around the elliptic fixed points. This suggests that bacterial activity enhances the strength of LVBs as transport barriers in chaotic flows. We also notice some differences in the size and shape of LVBs for the buffer and active cases. Since the LVBs are calculated from experimentally measured velocity fields, this indicates that microbial activity also modifies the underlying velocity field.
To quantify the strength of the transport barriers, or the reduction in dye transport, we calculate dye flux into the Lagrangian vortices using the following mass balance equation:
$\frac{\mathrm{d}}{\mathrm{d}t}\int\limits_{S(t)}C\,\mathrm{d}A=\oint\limits_{B(t)}D(\mathbf{\nabla}C\cdot\mathbf{n})\,\mathrm{d}l-\oint\limits_{B(t)}C(\mathbf{v}-\mathbf{v}_{B})\cdot\mathbf{n}\,\mathrm{d}l.$ (2)
Here, $S(t)$ and $B(t)$ denotes the enclosed area and the contour of the LVBs, $C$ is the dye concentration field, $D$ is the dye diffusivity, $\mathbf{n}$ is the normal vector of LVBs, $\mathbf{v}$ is fluid velocity, and $\mathbf{v}_{B}$ is the velocity induced by the motion and deformation of LVBs. The first and second terms on the right hand side of Eq. 2 are the contributions to the dye flux by diffusion and convection, respectively. The left hand side of Eq. 2 gives a simple way to calculate the dye flux as the rate of change of the total dye enclosed by the LVBs. Figure 2(c) shows the dye flux as a function of time, for vortex with label “4” in the buffer and the active suspension. We find a significant reduction of dye flux into the LVBs with bacterial activity, especially at earlier times ($N<150$). At later times ($N>200$), the dye flux in the active case is larger than that in the buffer, as the result of a larger diffusive flux due to large dye concentration gradients.
We now explore the generality of the reduction of dye flux in active suspensions. We start by calculating the circulation of a Lagrangian vortex, defined as:
$\Gamma(t)=\oint\limits_{B(t)}\mathbf{v}\cdot\,\mathrm{d}\mathbf{l}=\int\limits_{S(t)}\omega_{z}\,\mathrm{d}A.$ (3)
Here, $\omega_{z}$ denotes the vorticity component in the out-of-plane direction. We then define the the vortex Reynolds number as: $Re_{\Gamma}=\left|\overline{\Gamma}\right|/\nu$, where $\overline{\Gamma}$ denotes the time-averaged circulation over a period, and the absolute value is included since $\overline{\Gamma}$ can be positive or negative depending on the sign of the vorticity. We calculate $Re_{\Gamma}$ for each Lagrangian vortex in Figs. 2(a) and 2(b); the labels correspond to the magnitude of their $Re_{\Gamma}$. Inset of Fig. 2(c) shows the averaged dye flux as a function of the vortex Reynolds number $Re_{\Gamma}$, for the buffer and active cases. We find that in both cases the average dye fluxes scale linearly with $Re_{\Gamma}$. This is because both the average dye flux and $Re_{\Gamma}$ are linearly proportional to the area of the Lagrangian vortex. More importantly, however, we find that the slope of the scaling is larger in the buffer case compared to the active case. This result shows quantitatively that bacterial activity reduces the dye flux into the Lagrangian vortices and therefore enhances the strength of transport barriers.
3.2 Numerical Simulations
To understand how swimming microorganisms interact with an imposed time-periodic flow, we perform numerical simulations of swimming particles using the experimentally measured velocity field. Microorganisms are modelled as axisymmetric ellipsoids with a constant swimming speed $v_{s}$, in the direction $\mathbf{q}$ along their symmetry axis. The swimmer’s position $\mathbf{x}$ is modelled as:
$\dot{\mathbf{x}}=\mathbf{v}_{f}(\mathbf{x},t)+v_{s}\mathbf{q},$ (4)
where $\mathbf{v}_{f}$ is experimentally measured fluid velocity in the active suspension, and $v_{s}=20~{}\umu$m/s is the swimming speed of E. coli. As a control, simulations of elongated passive particles are also performed with $v_{s}=0$ in the velocity field measured in the buffer. The swimmer’s orientation is described by the Jeffery’s equation (Jeffery & Filon, 1922):
$\dot{\mathbf{q}}=[\mathbf{\Omega}(\mathbf{x},t)+\Lambda\mathbf{D}(\mathbf{x},t)]\mathbf{q}-\Lambda[\mathbf{q}\cdot\mathbf{D}(\mathbf{x},t)\mathbf{q}]\mathbf{q},$ (5)
where $\mathbf{D}$ and $\mathbf{\Omega}$ are the symmetric and skew-symmetric parts of the velocity gradient tensor, $\nabla\mathbf{v}_{f}$. Here, $\Lambda=(1-\alpha^{2})/(1+\alpha^{2})$ is a shape factor, with $\alpha$ being the swimmer aspect ratio; the factor is roughly $\Lambda\approx 0.88$ for E. coli.
Initially, both passive and active particles are uniformly distributed in the flow field with random orientations. As the simulations begin, passive and active particles start to develop complex patterns following the morphology of Lagrangian vortices (see Movie 3). Figures 3(a) and 3(b) show the spatial distribution of passive and active particles at $N=150$, respectively. The color code in the plots represents the local particle number density, $\rho_{N}$, normalized by the initial number density, $\rho_{0}$. We find that active particles are mostly depleted within the LVBs (green contours), suggesting that Lagrangian vortices repel elongated swimmers. This is in contrast to the passive case in which LVBs trap passive particles and lead to particle accumulation within the Lagrangian vortices. This results suggests the repulsion of active particles by the Lagrangian vortices is not a generic feature of the velocity field itself. Rather, it is the result of the interaction between Lagrangian vortices and elongated swimmers.
The repulsion of elongated swimmers by the LVBs is quantified by the radial distribution function: $g(r)=\langle\rho_{N}(r)\rangle/\rho_{0}$, where the angle bracket denotes a radial average in all directions calculated from the elliptic fixed points. The function $g(r)$ represents the likelihood of finding a particle at a radial distance $r$ from the elliptic fixed point of a Lagrangian vortex. Figure 3(c) shows $g(r)$ for both passive and active particles in the Lagrangian vortex with the label “4” in Fig. 2. The dashed line in Fig. 3(c) is a nominal radius of Lagrangian vortex, defined as: $r_{n}=\sqrt{A_{S}/\pi}$, where $A_{S}$ is the area enclosed by the LVBs. We find that the likelihood of finding an active particle within the nominal radius of a Lagrangian vortex is smaller than that of a passive particle, and vice versa for outside the nominal radius. This result further corroborates the repulsion of elongated active particles by Lagrangian vortices.
Figure 3: Spatial distributions of (a) passive particles, and (b) active particles, at time $N=150$. Particles are colored by their normalized local number density, $\rho_{N}/\rho_{0}$. Passive particles accumulate inside and active particles accumulate outside the LVBs (green contours). (c) Radial distribution function $g(r)$ calculated from the elliptic fixed points of the Lagrangian vortex with the label “4”; the dashed line is the nominal radius of the vortex. (d) Spatially averaged normalized number density within the LVBs, $\langle\rho_{N}\rangle_{S}/\rho_{0}$, as a function of time $N$ for the same vortex as in (c).
To quantify the time evolution of the repulsion of elongated swimmers, we calculate the average number density within the LVBs, $\langle\rho_{N}\rangle_{S}$, normalized by the initial number density $\rho_{0}$. Figure 3(d) shows $\langle\rho_{N}\rangle_{S}/\rho_{0}$ as a function of time $N$. Results show that $\langle\rho_{N}\rangle_{S}$ decreases non-monotonically with time for active particles, and drops to only 20% of the initial number density $\rho_{0}$ at $N=200$. This suggests that most active particles escape and accumulate outside the Lagrangian vortex as mixing progresses. In the control case, $\langle\rho_{N}\rangle_{S}$ for passive particles increases with time and reaches four times of the initial number density at $N=200$, which suggests the repulsion of elongated particles by LVBs is not an intrinsic feature of the velocity field. Although it has been observed that symmetric and circular Eulerian vortices can repel elongated swimmers (Torney & Neufeld, 2007; Ran et al., 2021; Qin & Arratia, 2022), here, we extend these results in a more general case of asymmetric and non-circular vortices from Lagrangian criteria.
Figure 4: (a) Stroboscopic trajectories of passive (blue) and active (red) particles that are initially inside the Lagrangian vortices illustrated by the TRA field. While sharing the same initial condition, passive particles remain trapped in the vortices and active particles spiral outward and escape. The trajectories of active particles are colored by their time, which appear more reddish as time progresses. (b) The PDFs of the inner product of the particle orientation vector, $\mathbf{q}$, and the tangent vector of members of elliptic LCSs, $\mathbf{t}$, as defined in the Inset. The initial condition ($N=0$) is approximately a uniform distribution for particles with random initial orientations. The PDFs at a later time ($N=50$) show that both passive and active particles preferentially align with members of elliptic LCSs at $\mathbf{q}\cdot\mathbf{t}=\pm 1$.
To gain further insights into how swimming particles are repelled by Lagrangian vortices, we plot the stroboscopic trajectories and orientations of passive and active particles as a function of time, as shown in Fig. 4(a). The Lagrangian vortices are illustrated by the TRA field (color map); passive and active particles are shown as blue and red bars, respectively. The trajectories of passive particles are blue, while active particles’ trajectories are colored by elapsed time that become more reddish with time. Despite sharing the same initial conditions for position and orientation, passive particles remain trapped in the Lagrangian vortices while active particles spiral outward with time and escape the vortices. We note that the swimming number of the active particles, $\Phi=v_{s}/U$, is at the order of $10^{-2}$ in the simulations, which suggests that the swimming speed is fairly negligible compared to the flow speed. However, even such (relatively) low swimming speed can drive active particles out of equilibrium and escape the Lagrangian vortices. That is, even small levels of swimming activity is enough to initiate the expulsion process. In addition, we find that the orientations of both passive and active particles tend to align with the level curves of the TRA field. Since the convex TRA contours locate the nested families of elliptic LCSs, this result indicates elongated particles—whether self-propelled or not—preferentially align with the members of elliptic LCSs. The difference is that the self-propulsion of active particles causes them to move outward and leave the vortices in a spiral manner.
Next, we test the alignment of the elongated particles, we calculate the inner product of the particle orientation vector, $\mathbf{q}$, and the tangent vector of the elliptic LCSs, $\mathbf{t}$. The inset of Fig. 4(b) shows a schematic of the two vectors. Since elliptic LCSs are nested families, we partition the elliptic LCSs into 10 members using the convex TRA contours. The inner product is calculated between the orientation vector of an elongated particle and the tangent of the nearest member of elliptic LCSs. Figure 4(b) shows probability density functions (PDF) of $\mathbf{q}\cdot\mathbf{t}$ for both passive and active particles at two different times. At $N=0$, the initial condition of the PDFs shared by both passive and active particles is approximately a uniform distribution due to the random initial orientations. At a later time of $N=50$, the PDFs become bimodal at $\mathbf{q}\cdot\mathbf{t}=\pm 1$, suggesting a parallel or anti-parallel alignment between $\mathbf{q}$ and $\mathbf{t}$; we do not find a substantial difference between the PDFs of passive and active particles, indicating self-propulsion does not play a significant role in affecting the alignment of elongated particles with elliptic LCSs. However, self-propulsion causes active particles to travel toward the outer members of the elliptic LCSs in both parallel and anti-parallel alignment configurations (see Movie 4), resulting in the spiral escaping trajectories of the active particles in Fig. 4(a). Consequently, the interplay between self-propulsion and the alignment of elongated particles with elliptic LCSs leads to the depletion of active particles inside the elliptic LCSs, that is, the accumulation of active particle outside the Lagrangian vortex boundaries (LVBs). This accumulation of microswimmers outside the LVBs can further obstruct dye transport into the Lagrangian vortices, which is responsible for the observed transport barriers.
4 Conclusion
In this manuscript, we investigate the interaction between swimming microorganisms and Lagrangian coherent vortices known as the elliptic LCSs in time-periodic flows in experiments and in simulations. Our results show that even small amounts of swimming activity can affect (i) the dynamics of active particles in the flow (Fig. 4) and consequently (ii) the mixing and transport of passive scalars (Fig 1) in chaotic flows. Experiments show that the interaction between organisms and elliptic LCSs leads to transport barriers through which the tracer flux is significantly reduced. Using the Poincaré map and the TRA field, we show that these transport barriers coincide with outermost member of elliptic LCSs, or Lagrangian vortex boundaries (LVBs). To further understand the formation of the transport barriers, we perform numerical simulations of elongated microswimmers in experimentally measured velocity fields. Results show that elliptic LCSs can repel elongated swimmer and lead to swimmer accumulation outside LVBs. This accumulation of microswimmers effectively reduces the transport into elliptic LCSs. We further show that the interplay between self-propulsion and the preferential alignment of elongated particle with the tangents of elliptic LCSs leads swimmers to escape the Lagrangian vortices. Overall, our results allow quantitative prediction of the Lagrangian transport of microorganisms and passive tracer quantities (e.g., temperature, oxygen, and nutrients) in chaotic flows with nontrivial vortex structures. Although there have been previous studies on the interaction between microorganisms and LCSs (Khurana & Ouellette, 2012; Dehkharghani et al., 2019; Ran et al., 2021; Qin & Arratia, 2022; Si & Fang, 2021, 2022; Yoest et al., 2022), our work extend study of swimmer-LCS interaction to elliptic LCSs (i.e., vortex-like dynamical structures). From a practical perspective, our results may be useful in understanding the organic matter flux in the oceans, algal blooms in lakes, and harmful bacterial infections.
Acknowledgement
We thank Tom Solomon, Kevin Mitchell, Simon Berman, and Albane Théry for insightful discussions, and Brendan Blackwell for help with early work. This work was supported by NSF-DMR-1709763.